OpenStudy (anonymous):

A survey of high school juniors found that 78% of students plan on attending college. If you pick three students at random, what is the probability that at least two plan on attending college? Round to the nearest percent. a) 53% b) 12% c) 40% d) 88%

OpenStudy (anonymous):

Please help!

OpenStudy (whpalmer4):

Well, for any 1 student you might select, what is the probability that they plan to attend college?

OpenStudy (anonymous):

50/50?

OpenStudy (whpalmer4):

How do you get that?

OpenStudy (anonymous):

well, there's only two choices and either one is a possibility.

OpenStudy (anonymous):

"at least two" means not none, not one compute \(P(x=0)\) and \(P(x=1)\) and subtract the result from 1

OpenStudy (anonymous):

oh i see, there are only three students so you could also compute \(P(x=2)\) and \(P(x=3)\) and add them

OpenStudy (anonymous):

binomial distribution \[P(x=2)=3\times (.78)^2\times (.22)\] and \[P(x=3)=(.78)^3\] if it is not clear where these numbers come from, let me know

OpenStudy (anonymous):

No, I understand. This is how I solved the equation but I got it wrong some how... \[3C2 (.78)^2 * (.22)^1\]

OpenStudy (anonymous):

Should I not have the 1 above .22?

OpenStudy (anonymous):

that is \(P(x=2)\) yes and \(_3C_2=3\)

OpenStudy (anonymous):

did you add \((.78)^3\) to your result?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

it says "at least two" so it means "two or three" you have to add them

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

So the answer's d-) 88% right?

OpenStudy (anonymous):

@satellite73 ?

OpenStudy (anonymous):

i don't know, i didn't compute it hold on i will check

OpenStudy (anonymous):

Thanks :)

OpenStudy (whpalmer4):

I get 0.876096 for that result..so yeah.

OpenStudy (anonymous):

I did too I just rounded. :)

OpenStudy (anonymous):

yeah, looks good

OpenStudy (anonymous):

Thanks so much for the help!